OpenStudy (anonymous):

Maggie's ma tells everyone that she is 14years younger than she actually is.if you add her actual age to the age that she tells people she is,the total is 122. What is the actual age of Maggie's ma(as reported by bob Dylan)?

OpenStudy (anonymous):

x = her age y = the age she tells people she is. x = y - 14 122 = x + y Substitute x in the equation 122 = x + y 122 = (y - 14) + y 2y - 14 = 122 2y = 136 y = 68 She is 86 years old.